Mr. Tanaka is 35years old and his son is 5, In how many years will Mr. Tanaka be three times as old as his son?
I have tried multiple times to figure out this problem, but I am sure I am doing the factors wrong..
F-fathers age
S-sons age
f*35=?
S=5
5x35=175
175/3= 59
so the answer that I am getting is 59, is this correct or am I doing this wrong?

Answer by sachi(548)

Mr. Tanaka is 35years old and his son is 5,
say after x years Mr. Tanaka be three times as old as his son
so after x yrs Mr. Tanaka will be 35+x yrs & his son will be 5+x yrs
so by question 35+x=3(x+5)=3x+15
or 3x-x=35-15
or 2x=20
or x=10
so after 10 years Mr. Tanaka be three times as old as his son